Construct a 99 ​% confidence interval about u1 - u2.
Population 1: n = 19, x = 17, s = 4.8
Population 2: n = 19, x = 21.2, s = 3.6

We are 99​% confident that the mean difference is between ... and ...

Answer by Boreal(15235)

t with df=36 (19+19-2), t=2.7195, difference in means is -4.2
s^2=sqrt((s1^2/19)+(s2^2/19))=1.89
s=1.38
99%CI half-interval is 2.72*1.38=3.75
CI=(-7.95, -0.45)